Given the points (-4,8) and (6, -12):

(i) Determine the midpoint of the line segment connecting the points.

(ii) Determine the distance separating the two points.


1
Expert's answer
2021-07-19T13:31:02-0400

Solution

"midpoint= \frac{X1+X2}{2}, \frac{Y1+Y2}{2}"

Midpoint for (-4,8) and (6, -12) ="\frac{-4+6}{2},\frac{8+-12}{2} =(1,2)"

Therefore the midpoint= "(1,2)"

Finding the distance separating the points;

"d=\sqrt{ (x2-x1)^2+(y2-y1)^2}"

"d= \sqrt{(10^2+(-20)^2}"

"d=\sqrt{500}"

"d = 22.36 units"